# MAPS: A Multilingual Benchmark for Agent Performance and Security

### v2

@@ 3.3 Translation Implementation and Verification
Translation Implementation Details. For NMT we used Google Cloud (2024). For LLM translation we used Cohere’s command A Cohere et al. (2025) and gpt-4o with a designated system prompt. Structured expressions such as Code, URL and equations were identified via the provided special tokens provided by the datasets Zhang et al. (2024); Hendrycks et al. (2021), and masked accordingly in the translation process.
Quality Assurance. To assess translation quality, we implemented a verification protocol involving native-speaking experts. Each sample was rated by a bilingual annotator fluent in English and the target language on a $1\textendash 5$ Likert scale across three established criteria: adequacy (semantic fidelity), fluency (grammatical and stylistic naturalness), and formatting accuracy (preservation of LaTeX, code, etc.) Freitag et al. (2021). A fourth metric, answerability, measured whether the translation preserved intent well enough for the annotator to confidently answer the question as if it were in English. Annotator instructions are provided in the Appendix. To validate verification reliability, we embedded “honeypot” samples with intentional errors; annotators reliably flagged these cases, confirming attentiveness and quality control. In total, 2,200 translations (25% of the benchmark) were verified, proportionally sampled across datasets and languages - for comparability, see Xu et al. (2025), where 1,600/6,000 tasks where verified.
Evaluation results confirm high translation quality across the benchmark, with an answerability rate of 94.294.2%, and mean scores of 4.43 (adequacy), 4.59 (fluency), and 4.75 (formatting) on a 1–5 Likert scale, supporting the benchmark’s preservation of semantic fidelity, linguistic naturalness, and structural integrity. Because a sample of n=2,000n=2{,}000 gives a worst-case 95% margin of error of $\pm 2.2$ percentage points (derivation in Appendix), the true corpus-level answerability lies in $92.0,\,96.4]\%$, indicating robust translation quality even after accounting for sampling uncertainty. Full per-language results and analysis are included in the Appendix. To support high-precision use cases, we also release a “verified” subset, consisting of 190~190 translations per language that passed native expert review across all four datasets.
@@ 4.1 Experimental Settings
Agent Assignment per Dataset. To demonstrate the utility of MAPS, we evaluate open-source agents on each multilingual dataset. We assigned a separate agent to each dataset, selected based on task relevance. While a unified agent would offer a more controlled evaluation, current systems lack the generalization needed to perform well across diverse tasks (Gioacchini et al., 2024; Chang et al., 2024). We employ each agent in its original configuration of prompts, system settings, tools and LLM backbone - all of which are multilingual models.
Figure 4: a) Multilingual effect as a function of the proportion of translated language tokens in input prompts. Each point represents a benchmark-agent pair, with the multilingual effect computed as the average relative degradation in performance or security across non-English languages. The trend suggests a correlation between input translation extent and multilingual vulnerability. b) Two snippets exemplify a low-translation prompt (MATH) and a high-translation prompt (GAIA), clarifying the x-axis percentages in panel (a) and showing how the proportion of natural-language tokens, rather than task difficulty alone, drives the observed multilingual effect.
@@
For GAIA, we used the OpenDeepResearch agent (von Platen et al., 2024), which integrates retrieval, web browsing, and tool use to support real-world reasoning. For MATH, we adopted MathChat (Wu et al., 2024), a zero-shot agent combining multi-turn reasoning with Python execution and the Wolfram Alpha tool. For SWE-bench, we applied SWE-agent (Yang et al., 2024b), which enables autonomous software reasoning through repository navigation, file editing, and test execution. For ASB, we built on the original benchmark infrastructure (Zhang et al., 2024), eliciting synthetic adversarial tool-use in various agentic scenarios (e.g. travel-agent, legal consultant, system admin).
We use OpenDeepResearch with GPT o1, MATHChat with GPT-4, SWE-agent with GPT-4.1, and ASB with Qwen2. Full configuration detailsare provided in the Appendix.
@@
Experiment Protocol. For each benchmark, the agent was evaluated three times in each of the $12$ target languages, yielding a total of $36$ runs per dataset. Fig. 3 reports the means and standard deviations over these runs.
Metrics. We adopt the original evaluation metrics from each benchmark to ensure consistency with prior agent evaluations. For MathChat (Math) and OpenDeepResearch (GAIA), we report the percentage of correct answers - with the latter matching either the English or translated reference. For SWE-agent, we report the percentage the percentage of submitted patches that successfully resolve the coding issue. For the ASB agent, we report the attack success rate (ASR). Additionally, we introduce a new metric: Multilingual Effect, which quantifies the performance or security gap between English and the average of all other languages. Given an evaluation metric $M$, the Multilingual effect is defined as follows:
$\text{Multilingual Effect }=\frac{1}{n}\sum_{i=1}^{n}M_{\text{lang}_{i}}-M_{\text{en}}$ (1)
Where $M_{en}$ denotes the performance in English, n is the number of non-English languages in the dataset (in our case $n=11$), and $M_{\text{lang}_{i}}$ represents the performance in the i-th non-English language.
@@ 4.2 Results
Fig. 3 presents the performance of open-source agents across all four benchmarks in English and the eleven target languages. In GAIA and ASB, we observe clear performance and security drops: non-English languages consistently underperformed compared to English, with reductions of up to 16% (GAIA) and a rise in vulnerability by up to 17% in ASB. Notably, SWE-bench and MATH exhibit only minor variation across languages, with most scores clustering around the English baseline. The Appendix extends this analysis with language-wise breakdowns of relative performance degradation.
These results reveal important differences in how multilingual degradation manifests across task types. Although all tasks require complex reasoning, some are more constrained than others. For instance, SWE-bench and MATH primarily involve structured code or mathematical expressions, placing less emphasis on natural language and more on Pythonic syntax or formal notation, thereby reducing the multilinguality effect. On the other hand, GAIA focuses on solving real-world tasks where understanding flexible natural language prompts is essential. Thus, the importance of the natural language problem statement is significantly higher. To explore this variation, we examine the proportion of localized, target-language text in each benchmark’s input (Fig. 4).
Interestingly, we observe that Chinese and Japanese yields the lowest ASR in the ASB benchmark, indicating the highest robustness to adversarial prompts in these two Asian languages. We attribute this pattern to the choice of backbone: the ASB agent was implemented using the Qwen2 model  (Yang et al., 2024a), which is known for its strong alignment for Chinese and Japanese language tasks. Qwen2 has consistently demonstrated strong performance in Chinese and Japanese-specific LLM benchmarks  [^3] https://rinnakk.github.io/research/benchmarks/lm/index.html, suggesting that alignment to a particular language and strategic backbone selection can significantly enhance resilience against multilingual adversarial prompts.
Fig. 4 examines the relationship between prompt composition and multilingual performance. part (a) shows a correlation between the percentage of non-English tokens in the input and the average performance gap (relative to English) across all four datasets. Benchmarks with higher proportions of localized, target-language-oriented input, such as GAIA and ASB, exhibit greater degradation, whereas SWE-bench, with predominantly English input (e.g., code), shows higher preservation. Part (b) demonstrate the differences between a MATH item with minimal translated text versus a GAIA query that is almost entirely translated.
@@
Orig. Trans.
GAIA (ACC) $\uparrow$ 47.4% 35.8% 41.3%
ASB (ASR) $\downarrow$ 48.8% 58.4% 56.6%
Table 1: Self-translation on GAIA and ASB. Making the agent aware of the multilingual context by self-translation of each non-English prompt recovers much of the lost performance/robustness, yet still leaves a noticeable gap to the native English baseline.
Self-translate Ablation. To further examine agents in multilingual settings, we followed the self-translate strategy of Etxaniz et al. (2024): each agent received one additional system message instructing it to translate any non-English prompt into English before solving the task, see results in Table 1. This led to improvements over direct multilingual inference, yet a clear gap to the English baseline persists, which may stem from the additional translation step introducing context loss or subtle distortions.
@@
Figure 5: Cross-lingual planning consistency for OpenDeepResearch and SWE-Agent. For each agent, we identify tasks where the model succeeds in English but fails in all other languages, then compute the semantic similarity between the English instruction and the initial planning step in each language using multilingual embeddings. OpenDeepResearch exhibits a strong cross-lingual alignment gap: in 85.7% of cases, the English planning step is more semantically faithful to the instruction than its counterparts in other languages. In contrast, SWE-Agent is more multilingual robust, with English outperforming other languages in only 25% of cases.
Cross-Lingual Failure Analysis: To investigate why some agents exhibit a more pronounced multilingual effect than others, we compare their initial task planning across languages—specifically, the agent’s first planning step, where it breaks down the instruction into a sequence of actions to solve the task. We compute the cosine similarity between the English instruction and each language’s corresponding planning step using multilingual embeddings from OpenAI’s text-embedding-3-small model OpenAI (2024b). This allows us to quantify how closely the agent’s reasoning in each language aligns with the original task intent expressed in English.
Our analysis focuses on cases where agents succeed in English but fail in all other languages. We compare two representative agents: OpenDeepResearchAgent on the GAIA dataset and SWE-Agent on the SWE-Bench dataset, as they exhibit the strongest and weakest multilingual effects, respectively (see Fig. 4).
As shown in Fig. 5, OpenDeepResearchAgent displays a strong English bias in planning alignment on GAIA: planning steps in English are more semantically aligned with the instruction than in other languages (0.76 vs. 0.70 on average), with lower variance. This indicates that the agent’s reasoning is more faithful to the instruction in English and degrades across languages. In contrast, SWE-Agent exhibits much less cross-lingual variation on SWE-Bench: planning similarity remains consistent across languages, suggesting that its reasoning process is less sensitive to linguistic differences.
These findings suggest that when an agent’s initial planning remains stable across languages, the impact of the multilingual effect is reduced. For instance, SWE-Agent follows a structured and language-invariant approach—reproducing and executing code—which helps maintain alignment regardless of language. Conversely, GAIA relies more on open-ended natural language reasoning, making it more susceptible to semantic drift across translations.
@@ 5.1 Guidelines for Multilingual Agent Deployment and Risk Assessment
Language-Composition-Aware Deployment Guidelines. Before deploying an AI agent in a multilingual setting, analyze the linguistic composition of its expected input, particularly the balance between structured elements (e.g., code, formal queries) and localized natural language. Inputs with a high proportion of natural language, tend to increase the risk of performance and safety degradation. We therefore recommend that in such case, developers should conduct a Multilingual Benchmark Assessment using a diverse, language-sensitive evaluation suite, such as ours, for AI agents operating across languages. This helps reveal vulnerabilities and promotes reliable real-world behavior across languages.
Multilingual Inputs Amplify Agentic Security Vulnerabilities. Our evaluation on ASB revealed that multilingual adversarial inputs can bypass agent safety mechanisms with minimal sophistication. Direct translations of English jailbreak prompts—without any adaptation or obfuscation—were sufficient to induce policy-violating behavior in multiple languages. This highlights a critical risk: localized inputs make simple attacks more effective, exposing security flaws in AI agents. Developers of safety-critical agentic systems should treat multilingual robustness as a core security concern and include translated prompts in safety evaluations using benchmarks like ours.
5.2 Benchmark Limitation
MAPS includes four datasets and eleven target languages, offering a strong foundation for multilingual evaluation. Nevertheless, extending these to additional domains, e.g. healthcare or legal reasoning, as well as to extremely low-resource languages (e.g., Amharic or Uyghur), would further enhance the benchmark. Nonetheless, the current suite already surfaces clear gaps across languages, offering valuable insights for multilingual deployment.

### v3

@@ 3.3 Translation Implementation and Verification
Translation Implementation Details. For neural machine translation, we used Google Translate Google Cloud (2024); for LLM translation, Cohere’s Command A Cohere et al. (2025) with a designated system prompt. Structured expressions (e.g., code, URLs, equations) were identified via special tokens provided by the datasets Zhang et al. (2024); Hendrycks et al. (2021), and masked accordingly in the translation process.
Figure 4: a) Multilingual Effect as a function of the proportion of translated language tokens in input prompts. Each point represents a benchmark-agent pair, with the Multilingual Effect computed as the average relative degradation in performance or security across non-English languages. The trend suggests a correlation between input translation extent and multilingual vulnerability. b) Two snippets exemplify a low‑translation prompt (MATH) and a high‑translation prompt (GAIA), clarifying the x‑axis percentages in panel (a) and showing how the proportion of natural‑language tokens, rather than task difficulty alone, drives the observed Multilingual Effect.
Quality Assurance. Translation quality was verified by native-speaking annotators fluent in both English and the target language. Each sample was rated on a 1–5 Likert scale across three established criteria: adequacy (semantic fidelity), fluency (grammatical and stylistic naturalness), and formatting accuracy (preservation of LaTeX, code, etc.) Freitag et al. (2021). A fourth metric, answerability, measured whether the translation preserved intent well enough for the annotator to confidently answer the question as if it were in English. Annotator instructions provided in the Appendix. To validate verifiers’ reliability, we embedded honeypot samples with intentional errors; annotators reliably flagged these cases, confirming quality control.
Evaluation results confirm high translation quality, with an answerability rate of 94.294.2%, and mean scores of $4.43$ (adequacy), $4.59$ (fluency), and $4.75$ (formatting) on the 1–5 Likert scale, demonstrating strong semantic fidelity, linguistic naturalness, and structural integrity. A total of 2,200 samples (25% of the benchmark) were manually verified, covering all languages and task types. At this verification scale, statistical analysis shows that even under the most conservative assumptions, the 95% confidence interval for the true corpus-level answerability rate is $92.0,\,96.4]\%$. This suggests that evaluating a quarter of the dataset provides a reasonably precise estimate of overall translation quality, with limited sampling uncertainty. Full per-language results and analysis are included in the Appendix. To support high-precision use cases, we also release a “verified” subset, consisting of approximately $190$ translations per language that passed native expert review.
Answerability as key metric. While adequacy and fluency confirm general translation quality, they do not ensure reliable task execution. Our correlation analysis shows that answerability, the extent to which a translation remains solvable, most closely aligns with correctness, whereas adequacy and fluency are not predictive. Although this relationship is not absolute since performance naturally also depends on task complexity, answerability remains the most reliable indicator of translation fidelity across languages (see Appendix for details).
@@ 4.1 Experimental Settings
Agent assignment per benchmark. To demonstrate the utility of MAPS, we evaluate one strong, established open-source agent per benchmark, each representing the current state of practice in its respective domain. While a single unified agent would offer a more controlled comparison, such general-purpose agents do not yet exist at sufficient maturity: contemporary agentic systems are tightly specialized around particular toolchains, environments, and prompting strategies, and transferring them across domains typically results in severe performance collapse (Gioacchini et al., 2024; Chang et al., 2024). Importantly, our goal is not to compare agent architectures, but to measure the effect of language on agentic behavior within a fixed agent–environment pair. For each benchmark, the agent, tools, prompts, and LLM backbone and configuration are held constant, and only the input language is varied. This isolates multilinguality as the sole experimental factor, ensuring that any observed degradation reflects language-conditioned failure modes rather than architectural differences.
@@
For GAIA, we employ the OpenDeepResearch (ODR) agent (von Platen et al., 2024), integrating retrieval, web browsing, and tool use for real-world reasoning. For MATH, we use MathChat (Wu et al., 2024), a zero-shot multi-turn reasoning agent with Python execution and the Wolfram Alpha tool. For SWE-Bench, we applied SWE-agent (Yang et al., 2024b) for autonomous code reasoning through repository navigation, file editing, and test execution. For ASB, we built on the original benchmark infrastructure (Zhang et al., 2024), eliciting synthetic adversarial tool-use in various agentic scenarios (e.g., travel, legal, system admin). ODR runs with GPT-o1, MathChat with GPT-4, SWE-Agent with GPT-4.1, and ASB with GPT-5-mini. Full configuration details are in the Appendix.
@@
Experiment Protocol. Each agent was evaluated three times per language, across all $12$ languages ($36$ runs per benchmark). We report mean and standard deviation over these runs in Fig. 3.
Metrics. We adopt each benchmark’s original evaluation metrics for consistency with prior work: accuracy for MATH and GAIA, test-case success rate for SWE-Bench, and attack success rate (ASR) for ASB. To enable cross-benchmark comparison of multilingual degradation, we further define a unified scalar measure, which we call the Multilingual Effect. For a given benchmark–agent pair, this is computed as the difference between the English performance score and the mean performance score across the 11 non-English languages. For ASB, where higher scores indicate worse security, the sign is inverted so that larger values consistently indicate greater degradation. Intuitively, the Multilingual Effect quantifies how much performance or safety is lost solely due to changing the interaction language. We use this measure throughout the analysis to study how linguistic load, task structure, and agent design modulate multilingual vulnerability.
@@ 4.2 Results
Fig. 3 presents agent performance across all four benchmarks in English and the eleven target languages. In GAIA and ASB, we observe a clear Multilingual Effect: non-English languages consistently underperformed compared to English, with reductions of up to 16% (GAIA) and a vulnerability increase of up to 27% in ASB. Notably, SWE-Bench and MATH exhibit only minor variation, with scores clustered around the English baseline.
These results reveal important differences in how multilingual degradation manifests across task types. Although all tasks require complex reasoning, some are more constrained than others. For instance, SWE-Bench and MATH primarily involve structured code or mathematical expressions, thus placing less emphasis on natural language and more on formal syntax or notation, thereby reducing the Multilingual Effect. In contrast, GAIA focuses on solving real-world tasks where understanding flexible natural language prompts is essential. Thus, the importance of the natural language problem statement is significantly higher. To explore this variation, we examine the proportion of localized, target-language text in each benchmark’s input (Fig. 4).
Fig. 4 examines the relationship between prompt composition and multilingual performance. Part (a) shows a correlation between the percentage of non-English tokens in the input and the average performance gap (relative to English). Benchmarks with higher proportions of localized, target-language-oriented input, such as GAIA and ASB, exhibit greater degradation, whereas SWE-Bench, with predominantly English input (e.g., code), shows higher preservation. Part (b) contrasts a minimally translated MATH sample example with a nearly fully translated GAIA query
@@
Orig. Determin. trans. Agent-dec. trans.
GAIA (ACC) $\uparrow$ 47.4% 35.8% 42% 40.9%
ASB (ASR) $\downarrow$ 32.7% 47.3% 40.4% 43.3%
Table 2: Comparison of deterministic and agent-decision translation on GAIA and ASB. We evaluate two ways to make the agent aware of multilingual context: (1) Deterministic translation, where inputs are pre-translated to English, and (2) Agent-decision translation, where the agent decides if and how to translate. Both recover part of the lost performance and robustness, though a clear gap to the English baseline remains.
Translation-based ablations. To disentangle translation effects from genuine multilingual degradation, we evaluate two strategies (Table 2). In deterministic translation, all non-English prompts are translated to English before inference, isolating translation quality effects. In agent-decision translation, the agent decides if and how to translate, reflecting realistic multilingual deployment. Deterministic translation recovers roughly 5% of the lost performance - consistent with our 5.8% human-verified translation error, while the multilingual gap remains higher (GAIA 12%, ASB 14%), indicating that translation bias alone cannot explain the degradation. The agent-decision variant yields comparable gains, showing partial self-correction through contextual awareness. Yet, both stay below the English baseline, pointing to deeper agent-level reasoning limits across languages.
Setting Benchmark(s) Configuration Multilingual Degradation (%)
Cross-Benchmark Check – Same Agent, Different Benchmarks
ODR MATH GPT-o1 –0.1%
MATHChat MATH GPT-4 –0.1%
ODR GAIA GPT-o1 –12%
Multiple LLM Check – Same Agent, Different Base LLMs
ODR GAIA GPT-o1, GPT-4o –12%,–10%
ASB Agent ASB GPT-5-mini, Qwen-2 –14.6%,–8.8%
Table 3: Cross-benchmark and cross-model analysis. Results indicate relative degradation compared to the English baseline. Multilingual performance remains stable in language-light tasks (MATH), but consistently degrades in language-heavy benchmarks (GAIA, ASB), independent of the underlying agent or base model.
Cross-Benchmark Evaluation. We examined whether multilingual degradation stems from agent designs or broader task characteristics. Such analyses are constrained by the benchmark-overfitting problem, where agents are tightly coupled to domain-specific tools and prompts, limiting cross-task generalization. We identified one viable case: using OpenDeepResearch (ODR), our GAIA agent, on the language-light MATH benchmark. ODR matched MathChat’s accuracy with no multilingual drop, yet degraded by 12% on GAIA, indicating that degradation arises primarily from linguistic load rather than agent architecture (see Table 3).
Cross-Model Evaluation. We next tested whether the multilingual degradation effect is specific to a particular LLM. Replacing GPT-o1 with GPT-4o for ODR and GPT-5-mini with Qwen-2 for ASB yields the same pattern: performance drops in GAIA and elevated ASR in ASB. These consistent trends suggest that multilingual degradation reflects a general property of language-conditioned reasoning, rather than a model-specific bias.
Unified-backbone check. A remaining concern is that multilingual differences across benchmarks might be caused by assigning different base LLMs to different agents (e.g., GPT-o1 for GAIA, GPT-4.1 for SWE-Agent, GPT-5-mini for ASB). While our primary setup preserves each agent’s native configuration to avoid unrepresentative failures (Section 4.1), we perform an additional experiment in which all agents are re-evaluated using the same backbone LLM, GPT-5-mini, while keeping all other components (agent logic, prompts, tools, and environments) fixed. Table 4 shows a re-evaluation of all benchmarks under this unified setting, which yields the same qualitative multilingual trends: GAIA and ASB continue to exhibit substantial multilingual degradation, whereas MATH and SWE-Bench remain comparatively stable across languages. Although absolute performance improves due to the stronger base model, the relative multilingual gaps persist, confirming that cross-benchmark differences are driven by task structure rather than model assignment. Full per-language results are provided in the Appendix.
Benchmark Original Unified
GAIA (Acc $\uparrow$) $12\%$ $10\%$
ASB (ASR $\downarrow$) $\% +14.6+14.6\%$
MATH (Acc $\uparrow$) $0.1\%$ $0.9\%$
SWE-Bench (Acc $\uparrow$) $1.3\%$ $0.3\%$
Table 4: Multilingual Effect (relative degradation vs. English) under the original agent setup and under a unified backbone (GPT-5-mini). Language-heavy benchmarks (GAIA, ASB) degrade substantially, while language-light benchmarks (MATH, SWE-Bench) remain stable.
@@
Figure 5: Cross-lingual planning consistency for OpenDeepResearch (ODR) and SWE-Agent. For each agent, we locate tasks solved in English but failed in all other languages, then measure the semantic similarity between the English instruction and each language’s initial planning step using multilingual embeddings. ODR exhibits a strong cross-lingual gap: in 85.7% of cases, its English planning step is more faithful to the instruction than in other languages. In contrast, SWE-Agent is more robust, with English leading in only 25% of cases.
Cross-Lingual Failure Analysis To identify the root causes of multilingual degradation, we conduct a deeper analysis of agents’ initial planning behavior across languages—the first stage where an instruction is decomposed into a sequence of actions. Using multilingual embeddings OpenAI (2024b), we measure the cosine similarity between each language’s plan and its English counterpart, quantifying how closely non-English reasoning aligns with the original task intent. We focus on tasks solved in English but failed in all other languages, comparing two contrasting agents: ODR (GAIA) and SWE-Agent (SWE-Bench)—which show the strongest and weakest Multilingual Effects, respectively (Fig. 4).
As shown in Fig. 5, ODR displays a clear English bias: its planning steps align more closely with task instructions in 85% of cases. In contrast, SWE-Agent shows minimal cross-lingual variation, indicating stable reasoning across languages. These results suggest that when an agent’s initial planning is language-invariant, multilingual degradation diminishes. A qualitative example of such an error is shown in the appendix.
@@ 5.1 Guidelines for Multilingual Agent Deployment and Risk Assessment
Guidelines for Language-Aware Deployment. Before deploying an AI agent in a multilingual setting, analyze the linguistic composition of its expected input, particularly the balance between structured elements (e.g., code, formal queries) and localized natural language. Inputs with a high proportion of natural language tend to increase the risk of performance and safety degradation. We therefore recommend that, in such a case, developers should conduct a Multilingual Benchmark Assessment using a diverse, language-sensitive evaluation suite, such as MAPS, for AI agents operating across languages. This helps reveal vulnerabilities and promotes reliable global real-world behavior.
Multilingual Inputs Amplify Agentic Security Vulnerabilities. Our ASB evaluation revealed that multilingual adversarial inputs can bypass agent safety mechanisms with minimal sophistication. Direct translations of English jailbreak prompts, without any obfuscation, were sufficient to induce policy-violating behavior across languages, exposing latent vulnerabilities in agent pipelines. Even advanced models such as GPT-5-mini proved more resilient in English than in translation. We thus recommend treating multilingual robustness as a security-critical dimension, and incorporating translated attacks as a standard component of agentic red-teaming risk assessments.
5.2 Agentic Robustness: Beyond LLM Multilingual Evaluation.
Our results show that multilingual robustness cannot be inferred from LLM-only benchmarks, since agentic systems introduce distinct failure modes absent from text generation. Even when an underlying LLM remains linguistically competent, agents can still fail when language misinterpretations propagate into operational errors such as planning drift or tool misuse (Fig. 5), reducing end-to-end reliability. Accordingly, multilingual robustness must be evaluated directly at the agent level, where performance and safety reflect language-conditioned decision making rather than text quality alone.